# Dynamic Contextual Pricing with Doubly Non-Parametric Random Utility Models

### v2

@@ Abstract
In the evolving landscape of digital commerce, adaptive dynamic pricing strategies are essential for gaining a competitive edge. This paper introduces novel doubly nonparametric random utility models that eschew traditional parametric assumptions used in estimating consumer demand’s mean utility function and noise distribution. Existing nonparametric methods like multi-scale Distributional Nearest Neighbors (DNN and TDNN), initially designed for offline regression, face challenges in dynamic online pricing due to design limitations, such as the indirect observability of utility-related variables and the absence of uniform convergence guarantees. We address these challenges with new population equations that facilitate nonparametric estimation within decision-making frameworks and establish the first analytical results on the uniform convergence rates of DNN and TDNN, enhancing their applicability in dynamic environments.
@@ 1 Introduction
Consumer demand is influenced not only by the price of a product but also by specific customer and product information, such as browsing history, zip codes, and product features. In digital sales environments, the economic model of random utility augmented with a wealth of sales data can offer crucial insights into consumer behavior, particularly consumers’ preferences and their reactions to different prices under varying contextual conditions. By implementing contextual pricing with random utility models, firms can not only gain deeper insights into the market, but also set dynamic prices tailored to individual customer profiles and product characteristics, potentially maximizing long-term revenue and gaining a competitive edge in the market.
Literature in dynamic contextual pricing often assumes a parametric form for either the mean utility, which represents the average preference of consumers, or the market noise, which captures the unobservable influences on demand. However, relying on parametric estimation poses significant disadvantages, primarily the risk of model misspecification. Parametric models make strong assumptions about the underlying data structure, and if these assumptions do not hold, the model’s predictions can be inaccurate, leading to suboptimal pricing strategies.
@@ 1.1 Doubly Nonparametric Random Utility Models
The proposed doubly nonparametric random utility model are characterized by (1) and (2) without any parametric characterizations on $\mathring{\mu}(\boldsymbol{x}_{t})$ or $F(\cdot)$. It operates within a broad economic framework of random utility models. In such settings, products are sold individually, with each sale resulting in a binary outcome indicating success or failure. At each decision point $t\in[T]$, a $d$-dimensional contextual covariate $\boldsymbol{x}_{t}$ from $\mathbb{R}^{d}$ is independently and identically drawn from a fixed, but a priori unknown, distribution. The expected market value of the product at time $t$ is determined by the observed contextual covariate $\boldsymbol{x}_{t}$. A general random utility model for the market value $v_{t}$ of the product is given by
@@
The doubly nonparametric contextual pricing works under a general economic framework of random utility models, the settings of which fundamentally differ from models that directly model the revenue function. For example, Chen and Gallego (2021) directly models the unknown expected revenue as $\operatorname{rev}_{t}(p_{t})=p_{t}\cdot d(\boldsymbol{x}_{t},p_{t})$ with $d(\boldsymbol{x}_{t},p_{t})\in[0,1]$ representing the personalized demand function. Bu et al. (2022) further assumes that $d(\boldsymbol{x}_{t},p_{t})$ separable in terms of $p_{t}$ and $\boldsymbol{x}_{t}$. In contrast, we work under the random utility model; the unknown expected revenue is given by (3) where it can be viewed as $d(\boldsymbol{x}_{t},p_{t})=1-F(p_{t}-\mathring{\mu}(\boldsymbol{x}_{t}))$. This finer granularity in demand decomposition allows us to separately analyze the mean utility function and noise distribution, thus improving theoretical guarantees. Furthermore, our analysis highlights a crucial trade-off between the dimensionality and smoothness of $\mathring{\mu}(\cdot)$ and $F(\cdot)$, offering practitioners a more nuanced control over pricing strategies.
@@ 1.2 Novelty and Contributions
Our theoretical developments include a new uniform convergence rate in Theorem 5 for the DNN and TDNN estimators. Employing the elegant U-Statistic representation of DNN and TDNN, Theorem 5 establishes that the DNN and TDNN estimators achieve the optimal rate of uniform convergence for nonparametric regression under the second-order smoothness condition and the fourth-order smoothness condition, respectively (Stone, 1982). Moreover, theoretical analyses uncover new insights into the interdependencies between the estimation accuracy for $\mathring{\mu}(\cdot)$ and $F(\cdot)$, and how these relationships influence the performance of the overall pricing strategy in the doubly nonparametric setting. Specifically, we demonstrate that the regret bounds for our doubly nonparametric dynamic pricing model depend intricately on the trade-offs between the dimensionality of contextual variables ($d$) and the smoothness properties of the noise distribution ($m$). To simply summarize, when $1\leq d\leq\frac{6}{m-1}$, the regret is bounded by $\widetilde{{\cal O}}(T^{\frac{2m+1}{4m-1}})$; and when $d>\frac{6}{m-1}$, the regret is bounded by $\widetilde{{\cal O}}(T^{\frac{d+4}{d+8}})$, ignoring the logarithmic terms. Regret bounds under both settings improve upon existing nonparametric models (Chen and Gallego, 2021; Bu et al., 2022), offering finer control and better performance in practical settings.
@@ 1.3 Literature and Organization
A stream of literature in contextual dynamic pricing models the demand function $d(\boldsymbol{x},p)$ or the revenue function $p\cdot d(\boldsymbol{x},p)$ directly (Qiang and Bayati, 2016; Ban and Keskin, 2021; Chen and Gallego, 2021; Wang et al., 2021; Bu et al., 2022). This stream of work can be categorized into contextual bandits with a specialized reward function $f(\boldsymbol{x}_{t},p)=p\cdot d(\boldsymbol{x},p)$ and continuous scalar action $p$. Under the general bandit setting, Slivkins (2011) studies nonparametric bandits assuming that $f(\boldsymbol{x}_{t},p)$ is Lipschitz and establishes the optimal regret being $\widetilde{{\cal O}}(T^{\frac{d+2}{d+3}})$ when $p$ is a scalar [^1] Slivkins (2011) works with a general setting where an action $p$ can be a vector.. Mao et al. (2018) exploit a special structure where the demand function is modeled as a strictly binary feedback $\mathbf{1}\{p\leq\mathring{\mu}(\boldsymbol{x}_{t})\}$, where $\mathring{\mu}(\boldsymbol{x}_{t})$ is Lipschitz, and develops the regret bound of $\widetilde{{\cal O}}(T^{\frac{d}{d+1}})$. Guan and Jiang (2018) apply $K$-Nearest Neighbour to nonparametric stochastic contextual bandits attaining a regret of $\widetilde{{\cal O}}(T^{\frac{d+1}{d+2}})$. Under the dynamic pricing setting, Chen and Gallego (2021) estimate $f(\boldsymbol{x}_{t},p)=p\cdot d(\boldsymbol{x},p)$ directly. However, they do not impose any functional structure that is specific to the dynamic pricing problem. Assuming that the revenue function is Lipschitz continuous and locally concave, the authors prove that the optimal regret is $\widetilde{{\cal O}}(T^{\frac{d+2}{d+4}})$ where $d$ is the dimension of $\boldsymbol{x}$. Nambiar et al. (2019) study the demand function $bp+g(x)$ motivated by model misspecification and propose a random price shock (RPS) algorithm to estimate price elasticity. They establish the optimal regret of order $\widetilde{{\cal O}}(\sqrt{T})$ compared to a clairvoyant who knows the best linear approximation to $g(x)$. Bu et al. (2022) further explore the separable structure of the demand function in the sense that $d(\boldsymbol{x},p)=a(\boldsymbol{x})+b(p)$ where both function $a(\cdot)$ and $b(\cdot)$ are unknown. This paper adopts the random utility model for demand characterization and is very different from these works in terms of aggregate demand modeling, algorithm development, and regret analysis.
@@ 2.1 Mean Utility Function Estimation
$\mathbb{E}[g_{t}\,|\,\boldsymbol{x}_{t}]$ $=\mathbb{E}[By_{t}\,|\,\boldsymbol{x}_{t}]=\mathbb{E}_{\varepsilon_{t}}\left[\mathbb{E}_{p_{t}}[By_{t}\,|\,\boldsymbol{x}_{t},\varepsilon_{t}]\;|\;\boldsymbol{x}_{t}\right]$ (5)
$=B\mathbb{E}_{\varepsilon_{t}}\Big{[}\frac{\mathring{\mu}(\boldsymbol{x}_{t})+\varepsilon_{t}}{B}\,\Big{|}\,\boldsymbol{x}_{t}\Big{]}=\mathring{\mu}(\boldsymbol{x}_{t}).$
@@ 2.3 Optimal Decision
Here $h\circ\mathring{\mu}(\cdot)$ is the composite function $h(\mathring{\mu}(\cdot))$ and $F^{(1)}$ is the first-order derivative of $F$.
@@ 3 Theoretical Guarantees
In this section, we establish the regret analysis of the proposed Algorithm 1 for doubly robust learning and decision in contextual dynamic pricing. For the theoretical analysis, we first introduce some notations and present necessary assumptions. For any candidate mean utility function $\mu:\mathcal{X}\mapsto\mathbb{R}$, let $\xi(z;\mu)$ be the density function of $Z=p_{t}-\mu(\boldsymbol{x}_{t})$ evaluated at $Z=z$ and define $F(z;{\mu}):=1-\mathbb{E}[y_{t}|p_{t}-{\mu}(\boldsymbol{x}_{t})=z]$. Therefore, the underlying true noise distribution $F(z)=F(z;\mathring{\mu})$ according to (11). We next introduce the assumptions about the market noise $\varepsilon_{t}$, the distribution function $F(z;\mu)$, density function $\xi(z;\mu)$, and the kernel function $K(x)$.
@@
There exist positive constants $L_{\phi}$ and $B_{u}$ such that $\phi^{(1)}(z)\geq L_{\phi}$ for all $z\in{\cal S}_{\varepsilon}$ with $\phi(z)$ defined in (17).
Assumptions 1, 2, and 3 are standard assumptions for the smoothness of the relevant functions in nonparametric estimation and inference. The rate of convergence $(T^{-\frac{2m+1}{(4m-1)(d+2)}}\land T^{-\frac{1}{2(d+4)}})\sqrt{d\log T}$ in Assumption 2 can be satisfied for the DNN estimator or TDNN estimator (with a further shrinked neighborhood size) with appropriately chosen subsampling sizes. Assumption 4 guarantees that $\phi(x)$ is strictly increasing and, thus, the uniqueness of the optimal solution of (16). It is weaker than that in Javanmard and Nazerzadeh (2019) since $1-F(u)$ is log-concave under the further restriction of $L_{\phi^{\prime}}\geq 1$.
The regret analysis unfolds in two steps. First, we will develop the uniform convergence rate of both the DNN and TDNN estimators in estimating the mean utility function $\mathring{\mu}$, as well as the convergence rate of the Nadaraya-Watson kernel estimator for estimating the noise distribution $F(\cdot)$. Second, building upon the uniform convergence rates established in the first step, we will derive an upper bound for the regret of Algorithm 1.
@@ 3.1 Statistical Convergence Rate of Learning
Suppose density function $f(\cdot)$ of $\boldsymbol{x}_{t}$ and the mean utility function $\mathring{\mu}(\cdot)$ are twice continuously differentiable with bounded second derivatives on the compact support $\mathcal{X}$ of $\boldsymbol{x}_{t}$. The density $f(\boldsymbol{x})$ is bounded away from 0 and $\infty$ for $\boldsymbol{x}\in\mathcal{X}$. Then we have for the DNN estimator that with probability $1-2\delta$,
@@
Assume DNN estimator $\widehat{\mu}_{k}^{\operatorname{DNN}}$ is applied with subsampling size $s=O(n_{k,\exp}^{\frac{d}{d+4}})$. Under condition of Theorem 5 for DNN estimator and Assumptions 14, there exist constants $C_{1}>0$ and $C_{2}>0$ such that for any episode $k$ satisfying $n_{k,exp}^{\frac{2m}{2m+1}}\geq\max(C_{1}d(\frac{\log n_{k,exp}}{2m+1}+1),3\log n_{k,exp})$ and $\delta\in(4\exp\{-C_{2}n_{k,exp}^{\frac{m}{2m+1}}/\log n_{k,exp}\},\frac{1}{2}]$, it holds with probability at least $1-2\delta$ that

### v3

@@ Abstract
In the evolving landscape of digital commerce, adaptive dynamic pricing strategies are essential for gaining a competitive edge. This paper introduces novel doubly nonparametric random utility models that eschew traditional parametric assumptions used in estimating consumer demand’s mean utility function and noise distribution. Existing nonparametric methods like multi-scale Distributional Nearest Neighbors (DNN and TDNN), initially designed for offline regression, face challenges in dynamic online pricing due to design limitations, such as the indirect observability of utility-related variables and the absence of uniform convergence guarantees. We address these challenges with innovative population equations that facilitate nonparametric estimation within decision-making frameworks and establish new analytical results on the uniform convergence rates of DNN and TDNN, enhancing their applicability in dynamic environments.
@@ 1 Introduction
The key in dynamic pricing is to model and estimate consumer demand function accurately. In practice, consumer demand is influenced not only by the price of a product but also by specific customer and product information, such as browsing history, zip codes, and product features. In digital sales environments, the economic model of random utility augmented with a wealth of sales data can offer crucial insights into consumer behavior, particularly consumers’ preferences and their reactions to different prices under varying contextual conditions. By implementing contextual pricing with random utility models, firms can not only gain deeper insights into the market, but also set dynamic prices tailored to individual customer profiles and product characteristics, potentially maximizing long-term revenue and gaining a competitive edge in the market.
Literature in dynamic contextual pricing often assumes a parametric form for either the mean utility (see, e.g., Shah et al. (2019); Fan et al. (2022); Luo et al. (2024)), which represents the average preference of consumers, or the market noise (see, e.g., Javanmard and Nazerzadeh (2019); Cohen et al. (2020); Xu and Wang (2021)), which captures the unobservable influences on demand. However, relying on parametric estimation poses significant disadvantages, primarily the risk of model misspecification. Parametric models make strong assumptions about the underlying data structure, and if these assumptions do not hold, the model’s predictions can be inaccurate, leading to suboptimal pricing strategies.
@@ 1.1 Doubly Nonparametric Random Utility Models
The proposed doubly nonparametric random utility model are characterized by (1) and (2) without any parametric characterizations on $\mathring{\mu}(\boldsymbol{x}_{t})$ or $F(\cdot)$. It operates within a broad economic framework of random utility models. In this paper, we consider the setting in which products are sold individually, with each sale resulting in a binary outcome indicating success or failure. At each decision point $t\in[T]$, a $d$-dimensional contextual covariate $\boldsymbol{x}_{t}$ from $\mathbb{R}^{d}$ is independently and identically drawn from a fixed, but a priori unknown, distribution. The expected market value of the product at time $t$ is determined by the observed contextual covariate $\boldsymbol{x}_{t}$. A general random utility model for the market value $v_{t}$ of the product is given by
@@
The doubly nonparametric contextual pricing operates within a general economic framework of random utility models, which diverge significantly from approaches that directly model the revenue function. For example, Chen and Gallego (2021) directly models the unknown expected revenue as $\operatorname{rev}_{t}(p_{t})=p_{t}\cdot d(\boldsymbol{x}_{t},p_{t})$ with $d(\boldsymbol{x}_{t},p_{t})\in[0,1]$ representing the personalized demand function. Bu et al. (2022) further assumes that $d(\boldsymbol{x}_{t},p_{t})$ is separable in terms of $p_{t}$ and $\boldsymbol{x}_{t}$. In contrast, we utilize the random utility model where the unknown expected revenue is given by (3) and the demand function is assumed to be $d(\boldsymbol{x}_{t},p_{t})=1-F(p_{t}-\mathring{\mu}(\boldsymbol{x}_{t}))$. The finer granularity in demand decomposition allows for separate analyses of the mean utility function and noise distribution, thus providing more precise control over the estimation errors of $F$ and $\mathring{\mu}$ and improving their statistical guarantees.
In addition, the structure of the random utility model facilitates an advanced analysis of how regret is influenced quadratically by the difference between optimal and estimated prices, which are closely tied to the estimation accuracy of $F$ and $\mathring{\mu}$. This connection leads to improved regret bounds under various scenarios. For further details on these relationships and their implications, refer to the discussions in Remark 3 in Section 3.2.
@@ 1.2 Main Contributions
Our theoretical developments include a new uniform convergence rate in Theorem 5 for the DNN and TDNN estimators. Employing the elegant U-Statistic representation (Lee, 2019) of DNN and TDNN, Theorem 5 establishes that the DNN and TDNN estimators achieve the minimax optimal rate of uniform convergence for nonparametric regression under the second-order smoothness condition and the fourth-order smoothness condition, respectively (Stone, 1982).
Moreover, our theoretical analyses reveal new insights into the interdependencies between the estimation accuracy for $\mathring{\mu}(\cdot)$ and $F(\cdot)$, and their impact on the overall performance of the pricing strategy within the doubly nonparametric setting. Firstly, our analysis shows that the regret bounds for our doubly nonparametric dynamic pricing model intricately depend on the trade-offs between the dimensionality of contextual variables ($d$) and the smoothness properties of the noise distribution ($m$), as detailed in the last row of Table 1. Secondly, the regret bounds for our methods outperform those of existing nonparametric models (Chen and Gallego, 2021; Bu et al., 2022), offering more precise control over varying smoothness levels of $F$ and $\mathring{\mu}$, and achieving enhanced performance in comparable scenarios.
Specifically, when the revenue function exhibits $2$nd-order smoothness, (i.e. $k_{a}=k_{b}=2$, $m=2$), our approach using the DNN estimator achieves a regret bound of $\widetilde{{\cal O}}(T^{\frac{5}{7}\lor\frac{d+4}{d+8}})$, , which is tighter than those achieved by other methods. Similarly, when the revenue function exhibits $4$th-order smoothness, (i.e. $k_{a}=k_{b}=4$, $m=4$), our method utilizing the TDNN estimator achieves the regret bound of $\widetilde{{\cal O}}(T^{\frac{3}{5}\lor\frac{d+8}{d+16}})$, again tighter than those offered by other models.
Paper Revenue Regret Key Assumptions
Slivkins (2011) $f(p,\boldsymbol{x})$ $\widetilde{{\cal O}}(T^{\frac{d+2}{d+3}})$ $f$: Lipschitz continuous
Chen and Gallego (2021) $f(p,\boldsymbol{x})$ $\widetilde{{\cal O}}(T^{\frac{d+2}{d+4}})$
$f$: Lipschitz continuous;
$f_{B}(p)$: smooth and concave
Bu et al. (2022) $p(a(\boldsymbol{x})+b(p))$ $\widetilde{{\cal O}}(T^{\frac{3}{4}\lor\frac{d+k_{a}}{d+2k_{a}}})$
$a$: $k_{a}$th-order smooth;
$b$: $k_{b}$th-order smooth
This work $p(1-F(p-\mathring{\mu}(\boldsymbol{x})))$ DNN: $\widetilde{{\cal O}}(T^{\frac{2m+1}{4m-1}\lor\frac{d+4}{d+8}})$
$F$: $m$th-order smooth;
$\mathring{\mu}$: 2nd-order smooth
TDNN: $\widetilde{{\cal O}}(T^{\frac{2m+1}{4m-1}\lor\frac{d+8}{d+16}\lor\frac{7}{13}})$
$F$: $m$th-order smooth;
$\mathring{\mu}$: $4$th-order smooth
Table 1: Comparison of the main results in this work and existing non-parametric dynamic pricing
@@ 1.3 Literature and Organization
A stream of literature in contextual dynamic pricing models the demand function $d(\boldsymbol{x},p)$ or the revenue function $p\cdot d(\boldsymbol{x},p)$ directly (Qiang and Bayati, 2016; Ban and Keskin, 2021; Chen and Gallego, 2021; Wang et al., 2021; Bu et al., 2022). This stream of work can be categorized into contextual bandits with a specialized reward function $f(\boldsymbol{x}_{t},p)=p\cdot d(\boldsymbol{x},p)$ and continuous scalar action $p$. Under the general bandit setting, Slivkins (2011) studies nonparametric bandits assuming that $f(\boldsymbol{x}_{t},p)$ is Lipschitz and establishes the optimal regret being $\widetilde{{\cal O}}(T^{\frac{d+2}{d+3}})$ when $p$ is a scalar [^1] Slivkins (2011) works with a general setting where an action $p$ can be a vector.. Mao et al. (2018) exploit a special structure where the demand function is modeled as a strictly binary feedback $\mathbf{1}\{p\leq\mathring{\mu}(\boldsymbol{x}_{t})\}$, where $\mathring{\mu}(\boldsymbol{x}_{t})$ is Lipschitz, and develops the regret bound of $\widetilde{{\cal O}}(T^{\frac{d}{d+1}})$. Guan and Jiang (2018) apply $K$-Nearest Neighbour to nonparametric stochastic contextual bandits attaining a regret of $\widetilde{{\cal O}}(T^{\frac{d+1}{d+2}})$. Under the dynamic pricing setting, Chen and Gallego (2021) estimate $f(\boldsymbol{x},p)=p\cdot d(\boldsymbol{x},p)$ directly. However, they do not impose any functional structure that is specific to the dynamic pricing problem. Assuming that the revenue function is Lipschitz continuous and locally concave, the authors prove that the optimal regret is $\widetilde{{\cal O}}(T^{\frac{d+2}{d+4}})$ where $d$ is the dimension of $\boldsymbol{x}$. Nambiar et al. (2019) study the demand function $bp+g(x)$ motivated by model misspecification and propose a random price shock (RPS) algorithm to estimate price elasticity. They establish the optimal regret of order $\widetilde{{\cal O}}(\sqrt{T})$ compared to a clairvoyant who knows the best linear approximation to $g(x)$. Bu et al. (2022) further explore the separable structure of the demand function in the sense that $d(\boldsymbol{x},p)=a(\boldsymbol{x})+b(p)$ where both function $a(\cdot)$ and $b(\cdot)$ are unknown. This paper adopts the random utility model for demand characterization and is very different from these works in terms of aggregate demand modeling, algorithm development, and regret analysis.
@@ 2.1 Mean Utility Function Estimation
$\mathbb{E}[g_{t}\,|\,\boldsymbol{x}_{t}]$ $=\mathbb{E}[By_{t}\,|\,\boldsymbol{x}_{t}]=\mathbb{E}_{\varepsilon_{t}}\left[\mathbb{E}_{p_{t}}[By_{t}\,|\,\boldsymbol{x}_{t},\varepsilon_{t}]\;|\;\boldsymbol{x}_{t}\right]=B\mathbb{E}_{\varepsilon_{t}}\Big{[}\frac{\mathring{\mu}(\boldsymbol{x}_{t})+\varepsilon_{t}}{B}\,\Big{|}\,\boldsymbol{x}_{t}\Big{]}=\mathring{\mu}(\boldsymbol{x}_{t}).$ (5)
@@ 2.3 Optimal Decision
Here $h\circ\mathring{\mu}(\cdot)$ denotes the composite function $h(\mathring{\mu}(\cdot))$ and $F^{(1)}$ is the first-order derivative of $F$.
@@ 3 Theoretical Guarantees
In this section, we establish the regret analysis of the proposed Algorithm 1 for doubly robust learning and decision in contextual dynamic pricing. For the theoretical analysis, we first introduce some notations and present necessary assumptions. For any candidate mean utility function $\mu:\mathcal{X}\mapsto\mathbb{R}$, let $\xi(z;\mu)$ be the density function of $Z=p_{t}-\mu(\boldsymbol{x}_{t})$ evaluated at $Z=z$ and define $F(z;{\mu}):=1-\mathbb{E}[y_{t}|p_{t}-{\mu}(\boldsymbol{x}_{t})=z]$. Therefore, the underlying true noise distribution $F(z)=F(z;\mathring{\mu})$ according to (11). We next introduce the assumptions about the market noise $\varepsilon_{t}$, the distribution function $F(u;\mu)$, density function $\xi(u;\mu)$, and the kernel function $K(x)$.
@@
There exists positive constants $L_{\phi}$ and $B_{u}$ such that $\phi^{(1)}(z)\geq L_{\phi}$ for all $z\in{\cal S}_{\varepsilon}$ with $\phi(z)$ defined in (17).
Assumptions 1, 2, and 3 are standard assumptions for the smoothness of the relevant functions in nonparametric estimation and inference. The rate of convergence $(T^{-\frac{2m+1}{(4m-1)(d+2)}}\land T^{-\frac{1}{2(d+4)}})\sqrt{d\log T}$ in Assumption 2 can be satisfied for the DNN or TDNN estimators (with a further improved neighborhood size) with appropriately chosen parameters. Assumption 4 guarantees that $\phi(x)$ is strictly increasing and, thus, the uniqueness of the optimal solution of (16). It is weaker than that in Javanmard and Nazerzadeh (2019) since $1-F(u)$ is log-concave under the further restriction of $L_{\phi^{\prime}}\geq 1$.
The regret analysis unfolds in two steps. Firstly, we will develop the uniform convergence rate of both the DNN and TDNN estimator in estimating the mean utility function $\mathring{\mu}$, as well as the convergence rate of the Nadaraya-Watson estimator for estimating the noise distribution $F(\cdot)$. Secondly, building upon the uniform convergence rate of the estimators, we will obtain an upper bound for the regret of Algorithm 1.
@@ 3.1 Statistical Convergence Rate of Learning
Suppose density function $f(\cdot)$ of $\boldsymbol{x}_{t}$ and the mean utility function $\mathring{\mu}(\cdot)$ are two times continuously differentiable with bounded second-order derivatives on the compact support $\mathcal{X}$ of $\boldsymbol{x}_{t}$. The density $f(\boldsymbol{x})$ is bounded away from 0 and $\infty$ for $\boldsymbol{x}\in\mathcal{X}$. Then we have for the DNN estimator that with probability $1-2\delta$,
@@
Assume DNN estimator $\widehat{\mu}_{k}^{\operatorname{DNN}}$ is applied with subsampling size $s=O(n_{k,\exp}^{\frac{d}{d+4}})$. Under condition of Theorem 5 for DNN estimator and Assumptions 14, there exist constants $C_{1}>0$ and $C_{2}>0$ such that for any episode $k$ satisfying $n_{k,exp}^{\frac{2m}{2m+1}}\geq\max(C_{1}d(\frac{\log n_{k,exp}}{2m+1}+1),3\log n_{k,exp})$ with probability at least $1-2\delta$ with $\delta\in(4\exp\{-C_{2}n_{k,exp}^{\frac{m}{2m+1}}/\log n_{k,exp}\},1/2]$, it holds that